# Object-level Geometric Structure Preserving for Natural Image Stitching

### v1

@@ 5 Experiments
Figure 3: APAP and SPHP see misalighment, and we delineated the indistinct portions using color-coded boxes. Autostitch, APAP, SPHP and GSP sees distortion. The convergence of the blue and red lines is essential, and we signify distortion by the intersection of these two lines. GES-GSP successfully prevents distortion, but it undergoes misalignment. Our method addresses misalignment and distortion well.
@@
In this section, we conduct a quantitative comparison among GSP [7], GES-GSP [13], and our proposed OBJ-GSP. Additionally, we qualitatively compared our results with those obtained from other image stitching methods, such as Autostitch [3], APAP [47] and SPHP [29], through visual assessment of the stitched images.
5.1 Experimental setup
@@
Evaluation metrics. We assess the quality of our stitching results from two perspectives. First, we employ the Mean Distorted Residuals (MDR) metric introduced in GES-GSP [13] to measure the degree of image distortion. In intuitive terms, if the points on the same side of the mesh were originally collinear and remain collinear after stitching, it implies that the stitching result has minimal distortion. Please refer to supplementary material of [13] for more details of MDR. Furthermore, we employ the Naturalness Image Quality Evaluator (NIQE) [33] metric to evaluate image quality and clarity. Compared to other image quality assessment methods like Structure Similarity Index Measure (SSIM) [41] and Peak Signal-to-Noise Ratio (PSNR), NIQE does not require a reference image. Since there is no strict ground truth for image stitching, we use NIQE to assess the quality of our stitching results. However, it is essential to emphasize that in the field of image stitching, there is no universally accepted objective evaluation method for assessing the quality of stitching results or the degree of distortion. Many practitioners still rely on subjective assessments based on human visual perception to gauge the quality of stitching outcomes.
@@
Datasets. We utilize datasets commonly associated with previous image stitching algorithms, which include datasets sourced from AANAP [28], APAP [47], CAVE [22], DFW [27], DHW [14], GES-GSP [13], LPC [20], SEAGULL [30], REW [26], SVA[32] and SPHP [29]. We also collected our own dataset of 18 image pairs for image stitching. Thus the datasets used in experiments encompass a diverse range of settings, including indoor and outdoor environments, various regions around the world, urban and natural landscapes, among others. Please kindly refer to the supplementary material for details on how we collected our own dataset and conducted an analysis of the ten datasets we use.
@@
Implementation details. We evaluated the results of GSP [7] and GES-GSP [13] using their publicly available C++ code. Our stitching code is also implemented in C++. Our implementation of SAM includes two approaches. Drawing from the Segment Anything C++ Wrapper [11], we exported the SAM [25] encoder and decoder into Open Neural Network Exchange (ONNX) format and subsequently replicated SAM’s automatic mode within C++. To achieve the best possible stitching results, we also directly implemented the semantic segmentation component using the SAM’s publicly available Python code and utilized their semantic segmentation results. The stitching part of our experiments ran on the CPU, while the SAM modules were capable of running on both CPU and GPU. Please kindly refer to the supplementary material on how we re-establish GSP and GES-GSP baselines.
5.2 Quantitative results
Table 1 shows MDR and NIQE results on datasets used in other stitching algorithms and our own dataset. The lower MDR and NIQE of OBJ-GSP indicates better stitching capability.
5.3 Qualitative results
@@
In the context of image stitching, human perception is primarily influenced by two key factors: alignment and distortion. Misalignment can result in undesirable artifacts such as ghosting and blurriness, while distortion can make the stitched image appear unnatural. Fig. 3 llustrates the stitching outcomes of six different methods, where we use straight lines and boxes to demonstrate the effects of alignment and distortion. GES-GSP addresses the issue of distortion in the stitching results, and building upon this foundation, our OBJ-GSP method achieves superior alignment performance. Fig. 1 and 4 elucidates the reasons behind the superior performance of our OBJ-GSP method. With the assistance of semantic segmentation techniques, we place greater emphasis on preserving critical structures and ensuring holistic protection at the object-level for objects within the images. Please kindly refer to our supplementary material for more results.
@@ 6 Ablation Studies
Semantic segmentation methods. To assess the influence of semantic segmentation results on the stitching performance, we conducted a comparative analysis across three different backbones of the SAM [25] model, namely ViT-B, ViT-L, and ViT-H [12], with a progression from smaller to larger models. Larger models are inherently capable of capturing more fine-grained semantic details. The corresponding stitching results are depicted in Figure 5. It is worth noting that the models based on ViT-B and ViT-L exhibit some blurriness and minor distortions. From the perspective of MDR, our model, under the three aforementioned backbone configurations, achieved improvements of 4.1%, 8.3%, and 9.4% over the baseline model GES-GSP [13], respectively.
@@
Figure 6: From top to down: one of the input images, OBJ-GSP with triangle samping and with triangle mesh sampling.
@@ 7 Discussion
Limitations. While our approach has demonstrated state-of-the-art performance in image stitching by extracting object-level geometric structures with SAM and preserving them with triangle mesh sampling, it still has several limitations:
1. Our method involves the use of large models for semantic segmentation, which results in longer processing times. According to our analysis, the effectiveness of geometric structure extraction significantly impacts the final results. Our method is constrained by the quality of SAM’s results. Smaller models like SAM [25] with Vit-B/L [12] do not perform as well as SAM with Vit-H. Fig. 5 demonstrates the differences.
2. As shown in Fig. 7, the proposed method fails in cases of significant parallax. In the perspective of the left image, the corner of the building appears to be obtuse. However, based on the inference from the right image, the corner of the building should be a right angle. Consequently, the stitching algorithm becomes perplexed, unsure of how to preserve the shape of the house. As discussed in [3], in an ideal scenario without parallax, one would only rotate the camera for capturing images without changing the position of the camera’s optical center. In general, we have not observed a significant improvement compared to GES-GSP in Fig. 7.
In only rare instances does OBJ-GSP experience failure due to the malfunction of SAM [25], such as in conditions characterized by inadequate illumination or an exceedingly sparse set of features. Please refer to the supplementary material for details.
@@
Seam-driven and deep learning based image stitching approaches. This paper primarily focuses on grid-based image stitching methods, which are currently the most explored and popular techniques. Previous work has explored seam-driven image stitching [15]. Seam-based methods, often combined with image blending [44, 45], have been widely used for post-processing in image stitching. The stitching approach discussed in this paper alleviates the need for extensive post-processing, resulting in more natural stitching results. As in provious grid-based works, we do not compare OBJ-GSP with seam-driven methods, but OBJ-GSP can be integrated with these post-processing methods. In recent years, several methods [21, 34] have attempted to model certain image stitching steps as unsupervised deep learning problems, leading to notable advances in the field.
Broader impacts. It’s important to acknowledge that our paper doesn’t explicitly discuss broader impacts in the proposed OBJ-GSP image stitching algorithm, such as fairness or bias. Segment Anything Model (SAM) [25] has discussed its broader impacts regarding geographic and income representation, as well as Fairness in segmenting people. Further research into how our algorithm may interact with other aspects of image stitching is encouraged.
@@ 8 Conclusion
In this paper, we propose Object-level geometric structure preserving for natural image stiching (OBJ-GSP) algorithm, which stands as a novel approach to achieving natural and visually pleasing composite images. The existing state-of-the-art, GES-GSP [13], while effective, faced limitations in accurately capturing geometric structures and preserving spatial relationships between these structures. In response, OBJ-GSP leverages semantic segmentation techniques, specifically the Segment Anything Model (SAM) [25], to extract object-level geometric structures imbued with semantic information. The proposed OBJ-GSP method addresses the shortcomings of its predecessor by preserving structures that align more intuitively with human perception. It employs triangular mesh deformations within individual objects to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets have unequivocally demonstrated that OBJ-GSP surpasses the performance of GES-GSP, establishing a new state-of-the-art benchmark in image stitching. In conclusion, the OBJ-GSP algorithm offers a promising direction for image stitching, enhancing the preservation of geometric structures and object-level details, ultimately leading to more natural and visually appealing composite images.
[…]
Supplementary Material
@@ 10 Reestablishing baselines
We replicated the results reported in the literature for GSP [7], GES-GSP [13], APAP [47], and SPHP [29] by implementing their publicly available codebases with their default parameter settings. For the structural alignment component, we employed the executable provided by Autostitch [3]. GES-GSP includes experimental data for both GES-GSP and GSP. In our experiments, the performance results for GES-GSP were consistent with the reported data, but GSP exhibited notably superior performance. This discrepancy may be attributed to variations in parameter configurations, as GES-GSP did not disclose the specific parameter settings used for the GSP experiments. In our method’s experiments, we maintained consistent parameter settings across all trials. Furthermore, for the structure extraction stage, we utilized the official code provided by SAM. However, we excluded masks with extremely small areas from our analysis.
@@ 11 More qualitative results
In this section we provide more qualivative results. We mark images with boxes to indicate misalignment, and use lines (and intersections of lines) to show distortion. Please refer to Fig. 8, 9, 10, 11 and 12. It is shown that our object-level preservation of structures can prevent distortion and misalignment at the same time.
@@ 13 Failure cases because of Segment Anything Model
We acknowledge that it’s very important to discuss the situations where, Segment Anything Model [25] fails, leading to the failure of the proposed OBJ-GSP image stitching algorithm. Please refer to Fig. 15 and 16. In the context of our approach, it is not imperative for the SAM [25] to achieve precise segmentation of objects containing semantic information. SAM need only recognize key object contours and boundaries. Therefore, the proposed OBJ-GSP is susceptible to SAM failure only in exceedingly rare instances where features are exceptionally sparse, and objects are highly indistinct. Consequently, SAM’s failure would impact solely the maintenance term of our structure, leading OBJ-GSP to degrade to the performance level of the GSP [7]. Simultaneously, we emphasize that SAM’s failure would result in the nullification of our structural preservation term, causing OBJ-GSP’s performance to regress to that of GSP only under scenarios where SAM proves ineffective. In addressing cases involving distortion and misalignment, we posit that mitigation strategies such as global straightening and multi-bend blending, as employed in Autostitch [3], can be leveraged to alleviate these issues.

### v2

@@ 5 Experiments
Figure 3: APAP [58] and SPHP [32] see misalignment, and we delineated the indistinct portions using color-coded boxes. Autostitch [3], APAP [58], SPHP [32] and GSP [7] sees distortion. The convergence of the blue and red lines is essential, and we signify distortion by the intersection of these two lines. GES-GSP [14] successfully prevents distortion, but it undergoes misalignment. Our method addresses misalignment and distortion well.
@@
Evaluation metrics. We quantitatively assess the quality of our stitching results from two perspectives: distortion prevention and alignment. First, we employ the Mean Distorted Residuals (MDR) metric introduced in GES-GSP [14] to measure the degree of image distortion. In intuitive terms, if the points on the same side of the mesh were originally collinear and remain collinear after stitching, it implies that the stitching result has minimal distortion. Please refer to supplementary material of [14] for more details of MDR. Furthermore, we employ the Naturalness Image Quality Evaluator (NIQE) [38] metric to evaluate alignment performance. NIQE measures image clarity, and stitching results with misalignment will produce blurry areas, leading to worse NIQE scores. Compared to other image quality assessment methods like Structure Similarity Index Measure (SSIM) [49] and Peak Signal-to-Noise Ratio (PSNR), NIQE does not require a reference image. Since there is no strict ground truth for image stitching, we use NIQE to assess the quality of our stitching results. However, it is essential to emphasize that in the field of image stitching, there is no universally accepted objective evaluation method for assessing the quality of stitching results or the degree of distortion. Many practitioners still rely on subjective assessments based on human visual perception to gauge the quality of stitching outcomes.
@@
Datasets. We utilize datasets commonly associated with previous image stitching algorithms, which include datasets sourced from AANAP [31], APAP [58], CAVE [25], DFW [30], DHW [15], GES-GSP [14], LPC [24], SEAGULL [33], REW [29], SVA[35] and SPHP [32]. We also collected our own dataset of 18 image pairs for image stitching. Thus the datasets used in experiments encompass a diverse range of settings, including indoor and outdoor environments, various regions around the world, urban and natural landscapes, among others.
Baselines. GSP [7] is a great work in alignment. Based on GSP, GES-GSP [14] protects line and curve structures. Our method tries to protect global structure based on GSP. The ideas behind these three methods are closely related. UDIS++ [40] is most famous as a pioneering work in deep learning based image stitching. Since it doesn’t explicitly use feature points, we are unable to measure its quality with MDR here. We provide a detailed comparasion between OBJ-GSP and UDIS++ in the supplementary materials.
@@
Implementation details. We evaluated the results of GSP [7] and GES-GSP [14] using their publicly available C++ code. Our stitching code is also implemented in C++. Our implementation of SAM includes two approaches. Drawing from the Segment Anything C++ Wrapper [12], we exported the SAM [28] encoder and decoder into Open Neural Network Exchange (ONNX) format and subsequently replicated SAM’s automatic mode within C++. To achieve the best possible stitching results, we also directly implemented the semantic segmentation component using the SAM’s publicly available Python code and utilized their semantic segmentation results. The stitching part of our experiments ran on the CPU, while the SAM modules were capable of running on both CPU and GPU.
Quantitative results. Table 1 shows MDR and NIQE results on datasets used in other stitching algorithms and our own dataset. We outperform GSP and GES-GSP in both alignment and shape preservation. UDIS++ [40] is a good try in deep learning based image stitching, but the performance is still no better than traditional methods(GSP, GES-GSP and ours).
@@
Qualitative results. In the context of image stitching, human perception is primarily influenced by two key factors: alignment and distortion. Misalignment can result in undesirable artifacts such as ghosting and blurriness, while distortion can make the stitched image appear unnatural. Fig. 3 illustrates the stitching outcomes of six different methods, where we use straight lines and boxes to demonstrate the effects of alignment and distortion. Fig. 1 and 4 elucidates the reasons behind the superior performance of our OBJ-GSP method. With the assistance of semantic segmentation techniques, we place greater emphasis on preserving critical structures and ensuring holistic protection at the object-level for objects within the images. Please kindly refer to our supplementary material for more quanlitative results.
@@ 6 Ablation Studies
Lightweight SAM backbones. To assess the influence of semantic segmentation results on the stitching performance, we conducted a comparative analysis across three different backbones of the SAM [28] model, namely ViT-B, ViT-L, and ViT-H [13], with a progression from smaller to larger models. Larger models are inherently capable of capturing more fine-grained semantic details. The corresponding stitching results are depicted in Fig. 5. It is worth noting that the models based on ViT-B and ViT-L exhibit some blurriness and minor distortions. From the perspective of MDR, our model, under the three aforementioned backbone configurations, achieved improvements of 1.9%, 2.7%, and 3.6% over the baseline model GES-GSP [14], respectively.
@@
Figure 6: From top to down: one of the input images, OBJ-GSP with triangle sampling and with triangle mesh sampling.
@@ 7 Discussion
Limitations: While OBJ-GSP has demonstrated state-of-the-art performance in image stitching by extracting object-level geometric structures with SAM and preserving them with triangle mesh sampling, it has several limitations:
1. Our method introduces a large semantic segmentation model into image stitching, resulting in higher computational costs. However, with the development of semantic segmentation techniques, lighter versions of SAM will emerge [53, 59], enhancing the speed of our work. According to our analysis, the effectiveness of geometric structure extraction significantly impacts the final results. Our method is constrained by the quality of SAM’s results. Smaller models like SAM [28] with Vit-B/L [13] do not perform as SAM with Vit-H. Fig. 5 demonstrates the differences. To stitch a pair of 800*600 images, SAM spends 25s on RTX2090 with 8-24G GPU memory, depending on the backbone. For C++ ONNX implementation, SAM VIT-B spends 1.5 min on CPU. Mesh optimization and image processing cost less than 4s on an Intel i5 CPU, almost the same as GES-GSP [14]. OBJ-GSP needs more computational resources and time than GES-GSP, but the stitching quality is also better. The time cost of SAM is larger than line detection methods in GES-GSP. The time used for triangle mesh optimization is almost the same as that in GES-GSP.
2. As shown in Fig. 7, the proposed method fails in cases of significant parallax. In the perspective of the left image, the corner of the building appears to be obtuse. However, based on the inference from the right image, the corner of the building should be a right angle. Consequently, the stitching algorithm becomes perplexed, unsure of how to preserve the shape of the house.
In only rare instances OBJ-GSP experiences failures due to the malfunction of SAM [28], such as in conditions characterized by inadequate illumination or an exceedingly sparse set of features. Please refer to the supplementary material for details.
@@
Applications of OBJ-GSP: In many fields, there is a need for high-quality stitched images, even at the expense of long time costs and significant computational resources. In medical image processing [46, 48, 61], for instance, stitching multiple pathological slice images to reconstruct the entire tissue structure or organ’s three-dimensional model demands high precision and quality. These tasks typically necessitate precise alignment and seamless fusion, and can tolerate longer computational times to ensure accuracy. Similarly, in photography [36] or virtual tourism [44, 50] applications, stitching numerous high-resolution images is necessary to generate high-quality panoramic images. This holds true in the fields of remote sensing [56, 9, 60] and movie industry [37, 19] as well.
Post-processing in image stitching: Our method primarily addresses computing transformation matrices to achieve alignment and shape preservation. Post-processing techniques can be combined with our approach to achieve a more natural stitching effect. Blending [3, 54, 55] and seam-driven [16] methods can be used to further reduce blurring, while global straightening [15] can decrease distortion.
Broader impacts: It is important to acknowledge that we do not explicitly discuss broader impacts in the proposed OBJ-GSP image stitching algorithm, such as fairness or bias. Segment Anything Model (SAM) [28] has discussed its broader impacts regarding geographic and income representation, as well as Fairness in segmenting people. Further research into how our algorithm may interact with other aspects of image stitching is encouraged.
@@ 8 Conclusion
In this paper, we propose Object-level geometric structure preserving for natural image stitching (OBJ-GSP) algorithm, which stands as a novel approach to achieving natural and visually pleasing composite images. The existing state-of-the-art, GES-GSP [14], while effective, faced limitations in accurately capturing geometric structures and preserving spatial relationships between these structures. In response, OBJ-GSP leverages semantic segmentation techniques, specifically the Segment Anything Model (SAM) [28], to extract object-level geometric structures imbued with semantic information. The proposed OBJ-GSP method addresses the shortcomings of its predecessor by preserving structures that align more intuitively with human perception. It employs triangular mesh deformations within individual objects to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets have unequivocally demonstrated that OBJ-GSP surpasses the performance of GES-GSP, establishing a new state-of-the-art benchmark in image stitching. In conclusion, the OBJ-GSP algorithm offers a promising direction for image stitching, enhancing the preservation of geometric structures and object-level details, ultimately leading to more natural and visually appealing composite images.
@@
9 Ablation of EfficientSAM and mesh sampling
We propose the utilization of SAM-based methods and mesh sampling to address distortion and misalignment during stitching. It is important to emphasize that both components are indispensable for object-level shape preservation. Fig. 8 illustrates the distinct results achieved under the same semantic segmentation output using line-based triangle sampling and object-level mesh sampling. Mesh sampling can recognize object structures and effectively preserve objects from distortion. Furthermore, with the advancement in the field of semantic segmentation, the speed of SAM-based methods is accelerating, which will greatly expedite our image stitching approach. For instance, EfficientSAM [53]. In Fig. 8, there is no significant difference observed between the results obtained using SAM + mesh sampling and EfficientSAM + mesh sampling. However, the time consumption of EfficientSAM is only $5\%$ of SAM, which is predictable. With the further development of SAM-based methods, even faster and more accurate approaches are expected to emerge, making our stitching method faster and more precise.
Figure 8: Ablation study of EfficientSAM [53] and mesh sampling.. The left images shows SAM [28] and EfficientSAM [53] segmentation results and sampling points on them. The purple boxes indicate misalignment in triangle sampling methods. Also, both triangle mesh based methods undergo distortion. We see no apparent difference between SAM [28]+mesh sampling and EfficientSAM+mesh sampling results.
@@ 11 Reestablishing baselines
We replicated the results reported in the literature for GSP [7], GES-GSP [14], APAP [58], and SPHP [32] by implementing their publicly available codebases with their default parameter settings. For the structural alignment component, we employed the executable provided by Autostitch [3]. GES-GSP includes experimental data for both GES-GSP and GSP. In our method’s experiments, we maintained consistent parameter settings across all trials. Furthermore, for the structure extraction stage, we utilized the official code provided by SAM. However, we excluded masks with extremely small areas.
@@ 12 More qualitative results
In this section we provide more qualivative results. We mark images with boxes to indicate misalignment, and use lines (and intersections of lines) to show distortion. Please refer to Fig. 9, 10, 11, 12 and 13. It is shown that our object-level preservation of structures can prevent distortion and misalignment at the same time.
@@ 14 Failure cases because of Segment Anything Model
We acknowledge that it’s very important to discuss the situations where, Segment Anything Model [28] fails, leading to the failure of the proposed OBJ-GSP image stitching algorithm. Please refer to Fig. 16 and 17. In the context of our approach, it is not imperative for the SAM [28] to achieve precise segmentation of objects containing semantic information. SAM need only recognize key object contours and boundaries. Therefore, the proposed OBJ-GSP is susceptible to SAM failure only in exceedingly rare instances where features are exceptionally sparse, and objects are highly indistinct. Consequently, SAM’s failure would impact solely the maintenance term of our structure, leading OBJ-GSP to degrade to the performance level of the GSP [7]. Simultaneously, we emphasize that SAM’s failure would result in the nullification of our structural preservation term, causing OBJ-GSP’s performance to regress to that of GSP only under scenarios where SAM proves ineffective. In addressing cases involving distortion and misalignment, we posit that mitigation strategies such as global straightening and multi-bend blending, as employed in Autostitch [3], can be leveraged to alleviate these issues.
15 Comparasion with UDIS++
UDIS [39](TIP 2021) and UDIS++ [40](ICCV 2023) are a family of attempts to address image stitching problems using deep learning frameworks. We compare with the revised version (UDIS++) as it performs better than UDIS. Like us, UDIS++ also aims to solve distortion issues on top of alignment. In the main text, we compute UDIS++’s NIQE to assess its alignment performance. Since UDIS++ is not feature point based, it cannot calculate MDR, a metric for measuring distortion. Therefore, in the supplementary material, we provide results for four scenes with multiple sets of images to intuitively compare distortion levels. Our method outperforms in both distortion resilience and alignment. Please refer to Fig. 18, 19 and 20.
@@ 15 Comparasion with UDIS++
Figure 18: Comparasion with UDIS++ on 6 sets of input images of a similar scene in its dataset. Line 1 and 3 are UDIS++ results, while line 2 and 4 are ours. Purple box indicates distortion of tree.
Figure 19: Comparasion with UDIS++ in its dataset. Purple box indicates distortion: the floor tiles are stretched.
Figure 20: Comparasion with UDIS++ in REW dataset [29]. UDIS++ stretches the stones and ground in purple box, while OBJ-GSP preserves the global structure.